# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v2

@@ 3.3 Local Existence Via Vorticity Formulation
We recall that in the general dimension $d$, IIE is shown to be locally well-posed in $H^{s+1}\times H^{s+1}$ for $s>d/2$, and we refer the readers to [Da06][[Da10]] [[DF11]] for various well-posedness results in $L^{p}$-based Sobolev spaces and endpoint Besov spaces. Here we show the local existence in Sobolev space of the vorticity formulation (47) of the inhomogeneous Euler equation.
We restrict ourselves to the torus case $D=\mathbb{T}^{2}$ for convenience. We’ll use the notation $H_{0}^{k}$ for space of mean-zero $H^{k}(D)$ functions, and we define the weighted space:
$\dot{H}_{\rho}^{1}:=\{f:\lVert\nabla f\rVert_{L_{\rho}^{2}}<\infty\}$
We start with the following elliptic estimate.
Lemma 3.3
Assume $\omega\in H_{0}^{2}$, $\rho\in H^{3}$ satisfying the assumption (A), then any mean-zero weak solution $\psi\in\dot{H}_{\rho}^{1}$ of $\mathbf{L}_{\rho}\psi=\omega$ also satisfies $\psi\in H_{0}^{4}$.
Proof: We’ll use a standard bootstrap argument. The proof is divided into $3$ steps.
Step 1.($H^{1}$ Estimate) Multiply both sides of the equation and integrating by part, we have:
$\lVert\nabla\psi\rVert_{L_{\rho}^{2}}^{2}=\int_{D}\lvert\nabla\psi(x)\rvert^{2}\rho(x)dx=-\int_{D}\psi(x)\cdot(\rho^{-1}\omega)(x)\rho(x)dx\leq\lVert\psi\rVert_{L_{\rho}^{2}}\lVert\rho^{-1}\omega\rVert_{L_{\rho}^{2}}.$
Now we control the $L_{\rho}^{2}$ norm of $\psi$ via $L_{\rho}^{2}$ norm of $\nabla\psi$ via Poincaré inequality. Thanks to assumption $A$ and Poincaré inequality, we have:
$\lVert\psi\rVert_{L_{\rho}^{2}}\leq C_{1}\lVert\psi\rVert_{L^{2}}\leq C_{1}C_{P}\lVert\nabla\psi\rVert_{2}\leq C_{1}^{2}C_{P}\lVert\nabla\psi\rVert_{L_{\rho}^{2}}.$
Where $C_{P}$ is the bounding constant of Poincaré inequality. We therefore have:
$C_{1}^{-2}\lVert\nabla\psi\rVert_{L^{2}}^{2}\leq\lVert\nabla\psi\rVert_{L_{\rho}^{2}}^{2}\leq C_{1}^{2}C_{P}\int_{D}\frac{\omega^{2}}{\rho}dx\leq C_{1}^{3}C_{P}\lVert\omega\rVert_{L^{2}}^{2}.$
Hence, we conclude $\lVert\nabla\psi\rVert_{2}\lesssim\lVert\omega\rVert_{2}$.
Step 2.($H^{2}$ Estimate) Now we expand the equation and normalize the leading coefficient to obtain:
$\Delta\psi+\nabla\psi\cdot\nabla\log\rho=\rho^{-1}\omega$ (50)
Simple rearrangement of (50) yields:
$\lVert\Delta\psi\rVert_{2}\leq\lVert\rho^{-1}\omega\rVert_{2}+\lVert\rho^{-1}\nabla\rho\cdot\nabla\psi\rVert_{2}\leq C_{1}\lVert\omega\rVert_{2}+C_{1}^{6}C_{P}\lVert\nabla\rho\rVert_{\infty}\lVert\omega\rVert_{2}.$
As a consequence of Sobolev embedding $H^{3}\hookrightarrow W^{1,\infty}$ in $2$-d, we have:
$\lVert\psi\rVert_{H^{2}}\lesssim(1+\lVert\rho\rVert_{H^{3}})\lVert\omega\rVert_{2}$
Step 3. (Bootstrap to Higher Order) Now we bootstrap the regularity in (50). A higher-order estimate yields:
$\lVert\Delta\psi\rVert_{H^{1}}\leq\lVert\rho^{-1}\omega\rVert_{H^{1}}+\lVert\rho^{-1}\nabla\rho\cdot\nabla\psi\rVert_{H^{1}}\lesssim\lVert\rho^{-1}\rVert_{H^{2}}\lVert\omega\rVert_{H^{1}}+\lVert\rho^{-1}\rVert_{H^{2}}\lVert\nabla\rho\rVert_{H^{2}}\lVert\psi\rVert_{H^{2}}$
where we use the algebra property of $H^{2}$ and the following para-product inequality (cf. [Wi22]):
$\lVert fg\rVert_{H^{1}}\leq\lVert f\rVert_{H^{1}}\lVert g\rVert_{H^{r}},\quad r>1$ (51)
Now, since $\rho\in H^{3}$ and the function $f(z)=z^{-1}$ is smooth on the range of $\rho$, we have $\rho^{-1}=f(\rho)\in H^{3}$. More precisely, we have
$\lVert\rho^{-1}\rVert_{H^{2}}=\sum_{j=0}^{2}\lVert D^{j}(1/\rho)\rVert_{2}\lesssim\sum_{j=0}^{2}\lVert\rho\rVert_{H^{2}}^{j}$
Hence, we conclude:
$\lVert\psi\rVert_{H^{3}}\lesssim\lVert\omega\rVert_{H^{1}}\lVert\rho^{-1}\rVert_{H^{2}}(1+\lVert\rho\rVert_{H^{3}}+\lVert\rho\rVert_{H^{3}}^{2})\lesssim\lVert\omega\rVert_{H^{1}}\sum_{j=1}^{4}\lVert\rho\rVert_{H^{3}}^{j}$
Finally, plugging the above estimate back to (50), we conclude
$\lVert\psi\rVert_{H^{4}}\lesssim\lVert\omega\rVert_{H^{2}}P(\lVert\rho\rVert_{H^{3}})$ (52)
where $P$ is some polynomial with positive coefficients. Hence, it is monotonically increasing on $(0,\infty)$. ∎
We now prove the local existence of IIE using our vorticity formulation, for simplicity we consider the casr $d=2$. Similar idea can be applied directly to the system (1) to conclude local well-posedness of strong solution in all dimensions. For a similar Lagrangian approach to the local well-posedness of hydrodynamic models, we refer the interested readers to [CI08][MP25][Zh10].
Theorem 3.4
@@
Assume $d=2$, Let $(\omega_{0},\rho_{0})\in H^{2}\times H^{3}$ and let $\rho$ satisfies Assumption (A), then (47) admits a unique solution local in time.
Proof. We consider the following iterative scheme:
@@
We aim to show that for sufficiently small time $T$ and positive numbers $R_{1},R_{2},R_{3}$ that we’ll choose later, the sequence $\{(\omega^{n},u^{n},\rho^{n})\}$ generated by the iterative scheme (1) stays in $U_{R}$.
@@
Now we consider the estimate for vorticity equation. By Duhamel’s principle, we can write down the following Lagrangian solution formulae:
@@
Hence, we may appeal to the algebra structure of $H^{2}$ and tandard result in propagation of Sobolev regularity along smooth flow (for instance, see [BCD11][MP25]) to estimate:
@@
where $C>1$ is some generic large number. Later we’ll fix $R_{2}$. Thanks to this choice, for all $t\in[0,T]$ we have:
@@
$\Delta(\psi^{n+1}-\psi^{n})+\nabla\log\rho^{n+1}\cdot\nabla(\psi^{n+1}-\psi^{n})+\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}=\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}+(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}$
Denote by $C_{R}$ constants depending on $R_{1},R_{2},R_{3}$ which may change from line to line. We now again make use of the (51) to control the RHS:
$\bigg\lVert\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\omega^{n+1}-\omega^{n}\rVert_{H^{1}},\quad\bigg\lVert(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}}$
and the commutator term
@@
Hence, following the same procedure in lemma 3.3, let $\delta\omega^{n}:=\omega^{n+1}-\omega^{n}$, $\delta\rho^{n}:=\rho^{n+1}-\rho^{n}$ and $\delta u^{n}=u^{n+1}-u^{n}$, we conclude
@@
$\partial_{t}\delta\omega^{n}+u^{n}\cdot\nabla\delta\omega^{n}+\delta u^{n}\cdot\nabla\omega^{n}=\frac{1}{2}(\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\}):=\sigma_{n}$
$\partial_{t}\delta\rho^{n}+u^{n}\cdot\nabla\delta\rho^{n}+\delta u^{n}\cdot\nabla\rho^{n}=0$
@@
$\delta\omega^{n}=\delta\omega_{0}\circ A_{t}^{n}+\bigg(\int_{0}^{t}(\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n}$
$\delta\rho^{n}=\delta\rho_{0}\circ A_{t}^{n}-\bigg(\int_{0}^{t}(\delta u^{n}\cdot\nabla\rho^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n}$
Follow the same line as for $\omega^{n}$, since $\delta\omega_{0}=\delta\rho_{0}=0$, we have the following Sobolev estimates:
$\lVert\delta\omega^{n}\rVert_{H^{1}}\leq e\int_{0}^{t}\lVert\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n}\rVert_{H^{1}}d\tau\leq e\int_{0}^{t}\lVert\sigma_{n}\rVert_{H^{1}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau$
$\lVert\delta\rho^{n}\rVert_{H^{2}}\leq e\int_{0}^{t}\lVert\delta u^{n}\cdot\nabla\rho^{n}\rVert_{H^{2}}d\tau\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau$ (56)
@@
$\{\lvert u^{n}\rvert^{2},\delta\rho^{n-1}\}+\nabla^{*}u^{n}\delta u^{n-1}\nabla^{\perp}\rho^{n-1}+\nabla^{*}\delta u^{n-1}u^{n-1}\nabla^{\perp}\rho^{n-1}$
@@
$\lVert\delta\omega^{n}\rVert_{H^{1}}\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n-1}\rVert_{H^{1}}+\lVert\delta\rho^{n-1}\rVert_{H^{2}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau$ (57)
Step 3. Now we define: $E^{n}(t):=\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{1}}$. By (55), (56), (57), we obtain:
@@
Therefore, we obtain $U_{n}\leq U_{n-1}/2\leq 2^{-(n-1)}U_{1}$. In particular, we demonstrate summability of sequence
$U_{n}=\sup_{t\in[0,T]}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{1}}.$
@@
Finally, closedness of $\mathcal{U}_{R}$ in $C([0,T];H^{1}\times H_{\sigma}^{2}\times H^{2})$ guarantees that $(\omega,u,\rho)\in\mathcal{U}_{R}$, which is the unique solution of (47). ∎
4 Lagrangian Analyticity of Solution
Lagrangian trajectories of many incompressible hydrodynamic models are known to be locally analytic in time [[FZ14]][CVW15], as long as the initial data lies in the local well-posedness class. On the other hand, such a nice property fails to hold when we switch to the compressible models (cf. [He19], since the pressure is no longer an elliptic constraint, but forces the system to be hyperbolic conservation law with finite speed of propagation. Therefore, there exists points which stays rest and then starts to move, prevents analyticity of Lagrangian displacement.
In this section, we approach the Lagrangian analyticity of inhomogeneous incompressible Euler equation using our Lagrangian formulation. Our approach is similar to [[FZ14]], where the vorticity formulation is used and Cauchy’s invariant plays a key role towards the iterative scheme for Taylor coefficients. In the present work, we stick to the momentum level, which allows us to deal with general dimensions in a unified way. The main theorem of this section is as follows.
Theorem 4.1
@@ 4 Lagrangian Analyticity of Solution
It turns out that inhomogeneity doesn’t prevent analyticity, by comparing above theorem 4.1 with the compressible Euler counterexample in [He19]. The crucial structure of incompressible fluids that leads to the analyticity is the hidden ellipticity in these models (The Biot-Savart law). Such a regularization mechanism, from the Lagrangian viewpoint, is manifested by the following pointwise integration by part lemma (cf. [Co01a][CI08]).
Lemma 4.2
@@
Proof. It suffices to show $\partial_{i}\mathbf{P}(\nabla^{*}\ell u)\in H^{s}$ for all $i$. For each fixed $1\leq i\leq d$, by commutativity between Fourier multipliers we have
@@
Proof of Theorem 4.1 We’ll use our Lagrangian representation to obtain a recursive formula for Taylor coefficients of Lagrangian displacement, and prove convergence of the Taylor series. The proof is divided into steps.
@@
Denote by $\ell_{t}(a)=X_{t}(a)-a$ the Lagrangian displacement, simplify the LHS, we conclude:
$\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\nabla(q_{t}\circ X_{t})+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert(\tau)d\tau=\eta_{0}$ (58)
@@
$\frac{d^{n}}{dt^{n}}\mathbf{P}\bigg[\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert^{2}(\tau)d\tau\bigg]=0,\quad\text{for all }n\geq 1$ (59)
@@
$\ell_{j}(t,a)=\sum_{k=1}^{\infty}t^{k}\ell_{j}^{(k)}(a)$
Notice here $\ell^{(0)}=X(0,a)-a=0$, $\ell^{(1)}=u_{0}(a)$. Straightforward computation yields:
$(\nabla^{*}\ell(t,a))_{ij}=\sum_{k=1}^{\infty}t^{k}\partial_{i}\ell_{j}^{(k)}(a),\quad\dot{\ell}_{j}(t,a)=\sum_{k=0}^{\infty}(k+1)t^{k}\ell_{j}^{(k+1)}$
Hence we moreover have:
$(\nabla^{*}\ell\dot{\ell})_{i}=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}(m-k+1)\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m+1-k)}$
$\int_{0}^{t}\lvert\dot{\ell}(\tau)\rvert^{2}d\tau=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}\frac{k(m+1-k)}{m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}$
@@
$\frac{d^{n}}{dt^{n}}\bigg\lvert_{t=0}\mathbf{P}\bigg[\rho_{0}\ell_{i}^{(1)}+\sum_{m=1}^{\infty}t^{m}\bigg((m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg)\bigg]=0$
@@
$\mathbf{P}\bigg[(m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg]=0$ (61)
Moreover determinant identity (60) naturally implies
@@
$\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\mathsf{Tr}\bigg[\bigg(\sum_{r=1}^{\infty}\nabla\ell^{(r)}t^{r}\bigg)^{k}\bigg]$
Hence the vanishing of coefficient for $t^{m}$ implies:
$0=\sum_{k=1}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})})$ (62)
@@
$U^{(m)}=\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)$

### v3

@@ 3.3 Local Existence Via Vorticity Formulation
We recall that in the general dimension $d$, IIE is shown to be locally well-posed in $H^{s+1}\times H^{s+1}$ for $s>d/2$, and we refer the readers to [14][15] [13] for various well-posedness results in $L^{p}$-based Sobolev spaces and endpoint Besov spaces.
Here, we show the local existence of strong solutions of IIE in Sobolev space via the vorticity formulation (42). We choose to work on a smooth and simply connected domain $D\subset\mathbb{R}^{2}$ for simplicity, but the proof of local well-posedness on $\mathbb{T}^{2}$ using (47) is also available following the same line of argument. We also expect that the Lagrangian–Eulerian system (25) yields a unified proof of local well-posedness for strong solutions in all dimensions. For a similar Lagrangian approach to the local well-posedness of hydrodynamic models, we refer the interested reader to [12][10][33][39]. However, we will only focus on the proof through our vorticity formulation to highlight its utility.
We start with the following elementary elliptic lemma. The proof is standard and is therefore omitted.
Lemma 3.5.
@@
Assume that $\omega\in H^{2}$ and $\rho\in H^{3}$ satisfies the assumption 2.1, then the solution $\psi$ of (43) satisfies the following estimate:
$\lVert\psi\rVert_{H^{4}}\lesssim P(\lVert\rho\rVert_{H^{3}})\lVert\omega\rVert_{H^{2}}.$ (48)
Here $P$ is some monotonically increasing polynomial function.
We now state our local existence result.
Theorem 3.6.
Assume $d=2$, Let $(\omega_{0},\rho_{0})\in H^{2}\times H^{3}$ and let $\rho$ satisfies the Assumption 2.1, then (42) admits a unique solution local in time.
Proof of Theorem 3.6.
We consider the following iterative scheme:
@@
The proof is Lagrangian in nature, we will use Sobolev estimate of transport equation to show existence of the fixed point of system (49). The proof is divided into $3$ steps.
@@
    We aim to show that for sufficiently small time $T$ and positive numbers $R_{1},R_{2},R_{3}$ that we will choose later, the sequence $\{(\omega^{n},u^{n},\rho^{n})\}$ generated by the iterative scheme (1) stays in $U_{R}$.
@@
    Now we consider the estimate for vorticity equation. By Duhamel’s principle, we can write down the following Lagrangian solution formula:
@@
    Hence, we make use of the algebra structure of $H^{2}$ and standard result in propagation of Sobolev regularity along smooth flow (for instance, see [4][33]) to estimate:
@@
    where $C>1$ is some generic large number. We leave $R_{2}$ to be fixed later. Thanks to this choice, for all $t\in[0,T]$ we have:
@@
    $\Delta(\psi^{n+1}-\psi^{n})+\nabla\log\rho^{n+1}\cdot\nabla(\psi^{n+1}-\psi^{n})+\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}=$
    $\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}+(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}.$
    Denote by $C_{R}$ constants depending on $R_{1},R_{2},R_{3}$ which may change from line to line. We now recall the following paraproduct inequality (see [38]): For $f\in H^{s}$, $g\in H^{r}$ with $r>s\geq 1$:
    $\lVert fg\rVert_{H^{s}}\leq\lVert f\rVert_{H^{s}}\lVert g\rVert_{H^{r}}.$ (51)
    By (51), we control the RHS as follows:
    $\bigg\lVert\frac{(\omega^{n+1}-\omega^{n})}{\rho^{n+1}}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\omega^{n+1}-\omega^{n}\rVert_{H^{1}},\quad\bigg\lVert(\frac{1}{\rho^{n+1}}-\frac{1}{\rho^{n}})\omega^{n}\bigg\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}},$
    and the commutator term satisfies
@@
    Hence, by standard elliptic estimate as in Lemma 3.5, let $\delta\omega^{n}:=\omega^{n+1}-\omega^{n}$, $\delta\rho^{n}:=\rho^{n+1}-\rho^{n}$ and $\delta u^{n}=u^{n+1}-u^{n}$, we conclude
@@
    $\partial_{t}\delta\omega^{n}+u^{n}\cdot\nabla\delta\omega^{n}+\delta u^{n}\cdot\nabla\omega^{n}=\frac{1}{2}(\{\lvert u^{n}\rvert^{2},\rho^{n}\}-\{\lvert u^{n-1}\rvert^{2},\rho^{n-1}\})=:\sigma_{n},$
    $\partial_{t}\delta\rho^{n}+u^{n}\cdot\nabla\delta\rho^{n}+\delta u^{n}\cdot\nabla\rho^{n}=0.$
@@
    $\delta\omega^{n}=\delta\omega_{0}\circ A_{t}^{n}+\bigg(\int_{0}^{t}(\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n},$
    $\delta\rho^{n}=\delta\rho_{0}\circ A_{t}^{n}-\bigg(\int_{0}^{t}(\delta u^{n}\cdot\nabla\rho^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n}.$
    Follow the same line of computation as for $\omega^{n}$, since $\delta\omega_{0}=\delta\rho_{0}=0$, we have the following Sobolev estimates:
    $\lVert\delta\omega^{n}\rVert_{H^{1}}\leq e\int_{0}^{t}\lVert\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n}\rVert_{H^{1}}d\tau\leq e\int_{0}^{t}\lVert\sigma_{n}\rVert_{H^{1}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau;$
    $\lVert\delta\rho^{n}\rVert_{H^{2}}\leq e\int_{0}^{t}\lVert\delta u^{n}\cdot\nabla\rho^{n}\rVert_{H^{2}}d\tau\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau.$ (53)
@@
    $\{\lvert u^{n}\rvert^{2},\delta\rho^{n-1}\}+\nabla^{*}u^{n}\delta u^{n-1}\nabla^{\perp}\rho^{n-1}+\nabla^{*}\delta u^{n-1}u^{n-1}\nabla^{\perp}\rho^{n-1}.$
@@
    $\lVert\delta\omega^{n}\rVert_{H^{1}}\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n-1}\rVert_{H^{1}}+\lVert\delta\rho^{n-1}\rVert_{H^{2}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau.$ (54)
  • Step 3. Now we define: $E^{n}(t):=\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}$. By (52), (53), (54), we obtain:
@@
    Therefore, we obtain $U_{n}\leq U_{n-1}/2\leq 2^{-(n-1)}U_{1}$. In particular, we demonstrate the summability of the sequence
    $U_{n}=\sup_{t\in[0,T]}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}.$
@@
    Finally, by our construction and thanks to Banach-Alaoglu theorem, the sequence $(\omega^{n},\rho^{n},u^{n})$ converges in weak-* sense in $\mathcal{U}_{R}$, whose limit must coincide with the strong $C([0,T];H^{1}\times H^{2}\times H^{2})$-limit $(\omega,u,\rho)\in\mathcal{U}_{R}$, which is the unique solution of (42).
4 Lagrangian Analyticity of Inhomogeneous Incompressible Fluids
Lagrangian trajectories of many incompressible hydrodynamic models are known to be locally analytic in time, as long as the initial data lies in the local well-posedness class. We refer the reader to [11][22][34][36][40] for detailed treatment on this subject. On the other hand, such a nice property fails to hold when we switch to the compressible models (see [23]), since the pressure is no longer an elliptic constraint, but forces the system to be hyperbolic conservation law with finite speed of propagation. Therefore, there may exist points which stay rest and then start to move, which prevents analyticity of the Lagrangian displacement. However, to the author’s knowledge, as an intermediate model between incompressible Euler and compressible Euler, whether such a property is enjoyed by IIE is not known in existing literature.
In this section, we approach the Lagrangian analyticity of the inhomogeneous incompressible Euler equation using our Lagrangian formulation. Our approach is similar to [40], where the vorticity formulation is used and Cauchy’s invariant plays a key role towards the iterative scheme for Taylor coefficients. In the present work, we stick to the momentum level, which allows us to deal with general dimensions in a unified way. The main theorem of this section is as follows.
Theorem 4.1.
@@ 4 Lagrangian Analyticity of Inhomogeneous Incompressible Fluids
It turns out that inhomogeneity doesn’t prevent analyticity, by comparing above theorem 4.1 with the compressible Euler counterexample in [23]. The crucial structure of incompressible fluids that leads to the analyticity is the hidden ellipticity in these models (The Biot-Savart law). Such a regularization mechanism, from the Lagrangian viewpoint, is manifested by the following pointwise integration by part lemma (cf. [12][10]).
Lemma 4.2.
@@
Proof of Lemma 4.2.
It suffices to show $\partial_{i}\mathbf{P}(\nabla^{*}\ell u)\in H^{s}$ for all $i$. For each fixed $1\leq i\leq d$, by commutativity between Fourier multipliers we have
@@
We now proceed to prove theorem 4.1.
Proof.
We will use our Lagrangian representation to obtain a recursive Taylor formula for the formal time series of Lagrangian displacement, and prove convergence of the formal time series. The proof is divided into $4$ steps.
@@
    Denote by $\ell_{t}(a):=X_{t}(a)-a$ the Lagrangian displacement, simplify the LHS, we end up with:
    $\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\nabla(q_{t}\circ X_{t})+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert(\tau)d\tau=\eta_{0},$ (55)
@@
    $\mathbf{P}\bigg[\rho_{0}(\nabla^{*}\ell+\mathbf{Id})\dot{\ell}+\frac{\nabla\rho_{0}}{2}\int_{0}^{t}\lvert\dot{\ell}\rvert^{2}(\tau)d\tau\bigg]=\mathbf{P}\eta_{0}.$ (56)
@@
    $\ell_{j}(t)=\sum_{k=1}^{\infty}t^{k}\ell_{j}^{(k)},$
    here $\ell_{j}$ is the $j$-th component of $\ell$. Notice that here $\ell^{(0)}=0$, $\ell^{(1)}=u_{0}$. Straightforward computation yields:
    $(\nabla^{*}\ell(t))_{ij}=\sum_{k=1}^{\infty}t^{k}\partial_{i}\ell_{j}^{(k)},\quad\dot{\ell}_{j}(t)=\sum_{k=0}^{\infty}(k+1)t^{k}\ell_{j}^{(k+1)}.$
    Hence, we moreover have:
    $(\nabla^{*}\ell\dot{\ell})_{i}=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}(m-k+1)\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m+1-k)};$
    $\int_{0}^{t}\lvert\dot{\ell}(\tau)\rvert^{2}d\tau=\sum_{m=1}^{\infty}t^{m}\sum_{k=1}^{m}\frac{k(m+1-k)}{m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}.$
@@
    $\mathbf{P}(\rho_{0}u_{0})=$
    $\mathbf{P}\bigg[\rho_{0}\ell_{i}^{(1)}+\sum_{m=1}^{\infty}t^{m}\bigg((m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg)\bigg],$
@@
    $\mathbf{P}\bigg[(m+1)\ell_{i}^{(m+1)}\rho_{0}+\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big)\bigg]=0.$ (58)
    Moreover, the determinant identity (57) naturally implies
@@
    $\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\mathsf{Tr}\bigg[\bigg(\sum_{r=1}^{\infty}\nabla\ell^{(r)}t^{r}\bigg)^{k}\bigg].$
    Hence, the vanishing of coefficient for $t^{m}$ implies:
    $0=\sum_{k=1}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m,r_{i}\geq 1}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})}).$ (59)
@@
    $U^{(m)}=\sum_{k=1}^{m}(m+1-k)\big(\partial_{i}\ell_{j}^{(k)}\ell_{j}^{(m-1+k)}\rho_{0}+\frac{k\partial_{i}\rho_{0}}{2m}\ell_{j}^{(m+1-k)}\ell_{j}^{(k)}\big).$